### v1

# Don’t be Fooled: The Misinformation Effect of Explanations in Human-AI Collaboration

Abstract.
Across various applications, humans increasingly use black-box artificial intelligence (AI) systems without insight into these systems’ reasoning. To counter this opacity, explainable AI (XAI) methods e enhanced transparency and interpretability. While recent studies have explored how XAI affects human-AI collaboration, few have examined the potential pitfalls caused by incorrect explanations. The implications for humans can be far-reaching but have not been explored extensively. To investigate this, we ran a study (n=160) on AI-assisted decision-making in which humans were supported by XAI. Our findings reveal a misinformation effect when incorrect explanations accompany correct AI advice with implications post-collaboration. This effect causes humans to infer flawed reasoning strategies, hindering task execution and demonstrating impaired procedural knowledge. Additionally, incorrect explanations compromise human-AI team-performance during collaboration. With our work, we contribute to HCI by providing empirical evidence for the negative consequences of incorrect explanations on humans post-collaboration and outlining guidelines for designers of AI.
1. Introduction
Imagine you are studying for an art history exam and must know how to distinguish two architectural styles. You seek advice from an online artificial intelligence (AI) assistant for explanations to differentiate the two styles. Despite being plausible, the AI’s explanation is incorrect. Yet you learn from these incorrect explanations, and as a consequence, your understanding of architectural styles is impaired! You fail your exam.
Recent rapid technological advancements, in particular generative AI with the popular example of ChatGPT, have significantly bolstered both the capabilities and the adoption of AI (Bommasani et al., 2021; Dwivedi et al., 2023). However, as AI technologies get more sophisticated, the complexity and opacity of supported decision-making processes increase as well, presenting new challenges to ensure that these systems are still transparent and interpretable for humans (Rudin et al., 2022). This increasing complexity necessitates a deeper understanding of the underlying factors impacting human-AI interaction, especially in scenarios where human oversight is critical (Amershi et al., 2019; Sterz et al., 2024). Regulatory frameworks, such as the EU AI Act, mandate human oversight to ensure AI systems operate ethically, legally, and safely (Commission, 2021). This underscores the significance of HCI research to develop methods and tools that facilitate effective collaboration between humans and AI (Shneiderman, 2020). Ensuring that AI is not only accurate but also explainable is vital for fostering appropriate reliance and complementary team performance (Schemmer et al., 2023c; Hemmer et al., 2023).
The need for explainability in AI has been widely acknowledged, leading to substantial advancements in both research and practice (Silva et al., 2023; Arrieta et al., 2020). In the financial sector, for instance, eXplainable AI (XAI) is used in credit scoring and fraud detection, targeting improved auditing, regulatory compliance, and user trust (Cirqueira et al., 2021).
Despite these advancements, the XAI literature has barely picked up on an important potential pitfall in human-AI collaboration scenarios: incorrect explanations (Morrison et al., 2024; Lakkaraju and Bastani, 2020) for accurate AI advice may impair humans post-collaboration as they compromise their procedural knowledge—their ability to perform specific tasks on their own— and their reasoning—their ability to derive conclusions based on their understanding. Exploring the effect of incorrect explanations on procedural knowledge and reasoning is crucial, as regulations (Commission, 2021) or performance expectations (Hemmer et al., 2023) might require humans to complement AI knowledge with their own knowledge and, therefore, maintain the ability to perform tasks on a superior performance level. This is particularly critical in high-stakes domains like healthcare, finance, and legal settings, where flawed decisions can have severe consequences (Rudin, 2019). Additionally, in scenarios where AI advice is first provided and then removed, like in AI-based teaching settings (Spitzer et al., 2023, 2024b), maintaining humans’ procedural knowledge is vital.
Understanding the impact of incorrect explanations in AI-assisted decision-making even when the AI advice is correct is essential for designing more effective collaborations between humans and AI. By identifying how and why incorrect explanations impact humans not only during, but in particular post-collaboration with AI, we can develop effective strategies to mitigate these effects (Chen et al., 2022; Lakkaraju and Bastani, 2020; Ehsan et al., 2019; Ribeiro et al., 2016). Thus, we ask the following research questions (RQs):
RQ1:
How do incorrect explanations for correct AI advice impair humans’ procedural knowledge in AI-assisted decision-making?
RQ2:
How do incorrect explanations for correct AI advice impair humans’ reasoning in AI-assisted decision-making?
RQ3:
How do incorrect explanations for correct AI advice affect the human-AI team performance?
To address these questions, we first synthesize how to measure the impact of incorrect explanations in AI-assisted decision-making based on previous research. We pre-registered our hypotheses on AsPredicted.org111We attached an anonymized copy of the pre-registration to the submission of this article.. Through an online study with 160 participants, we examine how incorrect explanations influence humans’ procedural knowledge (RQ 1) and reasoning (RQ 2) in a task to classify architectural styles of buildings. We provide participants with different types of AI support: no support at all, AI advice without explanations, AI advice with correct explanations, and AI advice with incorrect explanations. We measure the effects on their task performance during and after collaboration to derive the impact on their procedural knowledge and qualitatively analyze their understanding of the task to derive their reasoning ability. Additionally, we analyze the effect on the human-AI team performance (RQ 3).
The findings of our study reveal a misinformation effect in AI-assisted decision-making: incorrect explanations significantly impair humans, resulting in a notable decline in their procedural knowledge once they have to perform the task autonomously. We also find that humans’ reasoning is impaired when they receive incorrect explanations. In fact, a negative effect during the collaboration can also be substantiated in our study: human-AI teams perform worse when the AI provides incorrect explanations, curtailing the complementary benefits of this collaboration. These findings underscore the potential dangers of incorrect explanations and highlight the importance of developing robust and reliable explanatory support for humans.
In summary, our study makes several contributions to the field: first, it fills a critical gap in the literature by examining the impact of incorrect explanations on humans in AI-assisted decision-making. Second, it identifies and measures the misinformation effect within AI explanations and outlines its negative repercussions on humans—a decline in procedural knowledge and reasoning. Lastly, our research extends existing knowledge on the interplay between AI and human decision-makers, providing insights into the hazards of explanations on the human-AI team. Overall, this paper sheds light on the potential pitfalls of incorrect explanations and their implications for humans post-collaboration. With this exploratory work, we hope to contribute to the development of concepts and hypotheses helping to advance theoretical knowledge in XAI and AI-assisted decision-making as well as to further advance the design of more effective AI-based decision support systems.
2. Background
With the rapid advancement of AI and its integration into diverse decision-making processes, XAI has emerged as a critical technique for enhancing transparency and assistance to help decision-makers understand AI’s reasoning (Alufaisan et al., 2021). Especially with applications based on generative AI (like Open AI’s ChatGPT or Anthropic’s Claude), explanations are being generated in natural language that provide human-understandable support for the AI’s response (Singh et al., 2024; Zytek et al., 2024). Previous research on human-AI collaboration has predominantly focused on how AI explanations influence human decision-making (Subramanian et al., 2024; Schemmer et al., 2022). For example, Hemmer et al. (2021) examine the factors that impact human-AI team performance, suggesting that XAI can foster complementary collaboration between humans and AI. Albeit the positive effects, studies also highlight how XAI can impair the collaboration between humans and AI (Miller, 2019; Ehsan et al., 2021; Binns, 2018). However, the understanding of negative consequences of XAI is still limited (Mohseni et al., 2021), especially empirical evidence for the negative impact of incorrect explanations is missing. In Table 1, we sort recent works in AI-assisted decision-making according to the correctness of AI advice and explanations. The table shows the under-explored topic of incorrect explanations in HCI. We review recent literature that highlights the risks and limitations of collaborative settings between humans and AI, motivating the need for further exploration of how incorrect explanations impair humans.
The table shows in a two times two matrix how previous work contributed to correct/incorrect AI advice and correct/incorrect explanations. Correct AI explanations Incorrect AI explanations Correct AI advice (Schemmer et al., 2022; Hemmer et al., 2023; Schoeffer et al., 2024; Lai and Tan, 2019; Adhikari et al., 2019; Hase and Bansal, 2020; Ribeiro et al., 2018; van der Waa et al., 2021; Zhang et al., 2020; Yeung et al., 2020) (Morrison et al., 2024; Cabitza et al., 2024) Incorrect AI advice (Sadeghi et al., 2024; Schemmer et al., 2023c; Kim et al., 2023; Cau et al., 2023; Buçinca et al., 2021; Alufaisan et al., 2021; Ehrlich et al., 2011; Chen et al., 2023; Cabitza et al., 2023; Kocielnik et al., 2019) (Papenmeier et al., 2019; Morrison et al., 2024; Lakkaraju and Bastani, 2020)
2.1. Incorrect AI Advice in Human-AI Collaboration
Research in HCI has explored the effects of incorrect AI advice on human-AI collaboration (Kocielnik et al., 2019; Bansal et al., 2019). Kocielnik et al. (2019) propose three strategies for managing user expectations in scenarios involving incorrect AI advice: accuracy indicators, example-based explanations, and performance control. Through a study involving an AI-based scheduling assistant prone to errors, they demonstrate these techniques’ effectiveness in maintaining user satisfaction and acceptance despite occasional incorrect recommendations. Their findings also reveal that the specific nature of AI errors can significantly influence user perception and trust.
Several recent studies have investigated how programmers collaborate with Copilot, an AI programming assistant that is not always accurate (e.g., (Vasconcelos et al., 2023; Barke et al., 2023; Dakhel et al., 2023)). For instance, Vasconcelos et al. (2023) focus on an AI supporting code completion. As the AI can make errors in producing the code, they convey the uncertainty of Copilot’s outputs by highlighting code sections that are most likely to be edited by the programmer. Their findings reveal that this approach helps programmers to arrive at solutions more quickly. Additionally, Dakhel et al. (2023) conclude that, while GitHub Copilot is a valuable tool for expert programmers, non-expert programmers should exercise caution when using it due to the potential for errors. Further research in the HCI domain has delved into the broader implications of interacting with incorrect AI. Barke et al. (2023) examine how different levels of expertise among programmers influence their interaction with AI-generated code. Their results indicate that while expert programmers can effectively navigate and correct AI-generated errors, novice programmers often struggle, leading to decreased efficiency and increased frustration. Complementary studies have also explored how incorrect AI recommendations affect user trust and decision-making in various contexts. Yin et al. (2019) investigate user strategies for coping with incorrect AI advice in everyday tasks, discovering that users often develop heuristics to cross-verify AI recommendations with other information sources. Moreover, Yang et al. (2019) analyze the role of trust in human-AI collaboration, particularly in scenarios where AI occasionally provides incorrect advice. Their findings suggest that initial trust can facilitate smoother interactions, but that repeated exposure to incorrect advice significantly erodes trust.
2.2. Incorrect Explanations in Human-AI Collaboration
Next to the AI advice, the understanding in HCI of how incorrect explanations can have an impact in AI-assisted decision-making is limited. Only few studies investigate the effect of explanations’ incorrectness (Morrison et al., 2024; Cabitza et al., 2024; Lakkaraju and Bastani, 2020; Papenmeier et al., 2019). A recent study in HCI shows that not only incorrect advice but also incorrect explanations have the potential to deceive decision-makers (Morrison et al., 2024). The authors of Morrison et al. (2024) explore the negative impacts of such incorrect explanations on humans’ decision-making behavior. They extend the conceptualization of Schemmer et al. (2023c) by the explanation dimension and explore, in a bird classification study, how the correctness of explanations impacts humans’ reliance on AI. They show that incorrect explanations can deceive decision-makers, leading to an inappropriate reliance behavior. Cabitza et al. (2024) explore the effects of explanations in a logic puzzle task. They also show that incoherent explanations can mislead humans, resulting in an inappropriate reliance behavior. In another study, Papenmeier et al. (2019) run a study in AI-assisted decision-making with incorrect explanations. They study how the explanations affect humans’ trust in determining the publication of Tweets. Similarly, Lakkaraju and Bastani (2020) find that incorrect explanations can affect humans’ trust in AI by investigating their effects in law and criminal justice use-cases.
These studies collectively underscore the complex dynamics of human-AI interaction, especially how incorrect advice and incorrect explanations affect humans’ reliance behavior on AI. However, the HCI field lacks a deeper understanding of the effects of such impaired collaboration scenarios on humans themselves. Especially for scenarios in which not the advice but the explanation for the decision maker—intended to foster interpretability—is incorrect. Studies like Morrison et al. (2024) build a promising starting point to inform HCI researchers and practitioners of the downsides of incorrect explanations. However, we still do not know anything about the impact of these AI shortcomings on humans’ ability to perform the tasks autonomously (procedural knowledge) and to conclude about the underlying domain (reasoning) post collaboration.
3. Theoretical Development
In the evolving field of I, understanding the impact of AI explanations on decision-making has become critical. Explanations serve as a bridge between AI and humans, influencing trust, reliance, and collaboration (Schemmer et al., 2023c; Hemmer et al., 2023; Schoeffer et al., 2024). This work investigates how incorrect explanations, when paired with accurate AI advice, can mislead humans, potentially impairing their procedural knowledge and reasoning capabilities.
Schemmer et al. (2023c) highlight the role of AI advice in AI-assisted decision-making, emphasizing how they shape humans’ reliance on AI. Next to the advice, explanations influence the collaboration between humans and AI. Morrison et al. (2024) extend this understanding, exploring the effects of both correct and incorrect explanations on human reliance on AI. Their findings suggest that even when AI advice is accurate, incorrect explanations can distort human perceptions and decision-making processes, leading to an inappropriate reliance on AI. Building on these insights, our research focuses on a specific aspect of the work of Morrison et al. (2024): the misleading potential of incorrect explanations in the context of correct AI advice. While prior research has taken the first steps towards investigating the impact of incorrect explanations, still little is known about the repercussions on the ability of human recipients to perform tasks, namely their procedural knowledge, and to conclude based on their understanding, namely their reasoning, after the collaboration with the AI. We examine how such flawed explanations affect humans’ procedural knowledge, reasoning, and overall human-AI collaboration, contributing new insights to the growing body of literature on AI-assisted decision-making.
In AI-assisted decision-making, declarative knowledge (the ”know-what”) and procedural knowledge (the ”know-how”) are both crucial. Declarative knowledge is factual knowledge or information that humans hold. When exposed to incorrect explanations, incorrect explanations can distort this knowledge base, leading to flawed reasoning, as posited by Stenning and Van Lambalgen (2012).
Procedural knowledge refers to the ability to perform tasks and make decisions. The acquisition of procedural knowledge is highly sensitive to cognitive load, as discussed by Sweller (1988). Incorrect explanations can not only impair procedural knowledge, but also increase cognitive load, leading to difficulties in learning and applying procedures effectively. Chi and Wylie (2014) further support this by highlighting the importance of cognitive engagement in learning, which can be hindered by misinformation. Given these dynamics, we assume that incorrect explanations impair humans’ procedural knowledge. This is supported by Roediger and Butler (2011), who highlight how incorrect information encountered during learning can lead to the retention of false information, thereby impairing reasoning and memory. Similarly, Johnson and Seifert (1994) argue that incorrect information, even after correction, can still lead to wrong inferences. With prior research emphasizing the negative repercussions on procedural knowledge, we hypothesize:
Hypothesis 1:
Incorrect explanations lead to lower procedural knowledge.
When collaborating with AI, incorrect explanations may impair humans’ reasoning capabilities as they transition from knowing ”how” to understanding ”why”. In psychology, the work of Loftus et al. (1978) shows that misleading information can impair human memory, leading to an erroneous understanding. Ecker et al. (2011) demonstrate the persistence of misinformation effects, showing how they continue to influence reasoning even after correction, thus highlighting their lasting impact. Similarly, Kendeou et al. (2013) investigate the effects of incorrect information on human reasoning and explore mechanisms to mitigate these effects. Building upon prior research that examines how misinformation impairs decision-making and reasoning (Soon and Goh, 2018), we adopt this perspective to AI-assisted decision-making and hypothesize the following:
Hypothesis 2:
The reasoning for the classification in the task is impaired by incorrect explanations.
Seeber et al. (2020) emphasize the importance of communication for effective collaboration between humans and AI. They argue that misunderstandings or unclear communication can lead to disruptions in team performance, particularly when AI systems provide incorrect or unclear explanations. Furthermore, Dzindolet et al. (2003) investigate how trust in automated systems is impacted by incorrect feedback, leading to reduced team performance. Similarly, Eiband et al. (2018) provide insights into how human-AI collaboration is affected by the transparency of AI systems, suggesting that incorrect advice can hinder effective teamwork. Lastly, Bansal et al. (2021) explore the dynamics of human-AI interaction, showing that incorrect AI advice alongside explanations can create friction in collaboration, leading to poorer outcomes. Incorrect explanations in AI-assisted decision-making can impair the collaboration, leading to inappropriate reliance on AI (Morrison et al., 2024). With prior research showing the relationship between humans’ reliance behavior on AI and the human-AI team performance (Schemmer et al., 2023c), we assume that the human-AI team performance drops when humans are provided with incorrect explanations:
Hypothesis 3:
The collaboration with a misleading AI leads to a lower human-AI team performance.
4. Methodology
In this section, we describe our methodology to assess how incorrect explanations for correct AI advice influence humans during and post collaboration with an AI. We set up an online study and investigated how participants performed in a visual classification task on an architectural dataset. In this section, we outline the task domain, the study design, the recruitment of participants, the development of the AI, and finally, the metrics that we use to assess our RQs. Before we ran the study, we pre-registered the study on AsPredicted.org to report our hypotheses, our treatments, our planned analyses, and our exclusion strategy. An anonymized copy of the pre-registration is provided in the supplemental materials.
4.1. Task Domain
In order to analyze the impact of explanations, we chose to use a task that most people are not familiar with and typically cannot handle themselves initially: the classification of the architectural style of buildings. To do so, we use the established dataset of Xu et al. (2014) containing images of buildings across 25 different architectural styles. In close discussion with architecture researchers of the local university, we chose three architectural styles that share similar features and are not easily distinguishable: Art Nouveau, Art Deco, and Georgian Architecture. For each architectural style, we selected 30 images that clearly represent the features of each architectural style and, thus, are appropriate instances for our study. We further made sure that the buildings were centered in each image and cropped all irrelevant information in the images, like other buildings.
4.2. Study Design
Our research questions target the understanding of the impacts of incorrect explanations in AI-assisted decision-making on human procedural knowledge and reasoning and the resulting human-AI team performance. To address them, we employed a study combining between- and within-subjects design: between subjects, we analyze the impact of different AI support. Within subjects, we observe this impact in different stages of decision-making: before, during, and after collaboration with the AI. The study was approved by the university’s institutional review board.
The figure shows the study procedure divided into five parts: The introduction, the pre-test, the main task, the post-test and the questionnaire.
The online study was divided into five different parts (see Figure 1): i part (1), the participants had to give their consent to participate and were introduced to the study and its procedure. They also received context information about the three different architectural styles with a description of their main characteristics (see Table 3 in Appendix A). Additionally, we included two attention checks, one of them in part (1): we asked participants what their task would be throughout the study, and they could select from three options. In part (2), a pre-test assessed participants’ task performance in classifying architectural styles of buildings: they were each shown six images randomly drawn from a bucket of 18 pre-selected images of the dataset. We ensured that each class was balanced for each participant (two images per class) and that the order in which the classes were shown to participants was varied. By doing so, we minimized the risk of inducing biases in the order of images. It also ensured that our results were not dependent on the difficulty of pre-selected images. Moreover, the pre-test allowed us to check that participants in each treatment did not differ in their prior expertise regarding the task. Following the pre-test, participants were randomly assigned to a treatment in the main task in part (3), distinguishing the type of AI support received:
Control group: Participants did not receive any AI support
AI treatment: Participants received only the AI’s classification
Correct explanation treatment: Participants received the AI’s classification and a correct explanation
Incorrect explanation treatment: Participants received the AI’s classification and an incorrect explanation
By distinguishing the groups in this way, we can infer the effect of incorrect explanations on participants’ procedural knowledge and reasoning. After the assignment, participants had to classify buildings’ architectural styles on twelve images (see as an example Figure 2(b)). All twelve images were randomly drawn from a bucket of 30 pre-selected images balanced in classes. Similar to the pre-test, each participant was provided with four images of each class, and the order was randomized. Each image was displayed on a separate page in the study. During the main task, participants had the option to click on a button to show the context information for the architectures’ characteristics and verify the support they received. In part (4), the post-test, participants of each treatment had to classify six different images without support. Similar to the pre-test, the images were drawn from another bucket of 18 pre-selected images balanced in classes. Each participant was shown two images of each class on separate pages in randomized order.
To prevent participants from saving time by just randomly clicking through the study, participants could continue to the next page in the pre-test, main task, and post-test only after a few seconds. This design choice followed the protocol of (Spitzer et al., 2024a) and was to ensure that participants focused on the actual task. In the final part (5), participants answered a questionnaire. The questionnaire asked them to describe and explain each architectural style. With this, we were able to analyze whether participants had learned the correct distinctions between the architectural styles and were able to conclude based on their understanding (e.g., their reasoning). On top of that, they had to rate several variables on seven-point Likert scales to measure for confounding factors. These included the AI usefulness, their experience with AI, their cognitive load, and their trust in AI. All variables are presented in Appendix A in Table 5. Throughout the questionnaire, we implemented a second attention check to ensure only valid results, as suggested by Abbey and Meloy (2017). In this attention check, we asked participants to select ”Likey“ for an item of AI usability (”As an attention check, please choose ”Likely” for this statement“). The questionnaire ended with demographic questions on participants’ age, gender, education and employment.
The figure shows two examples of instances of the main task in the correct explanation treatment on the left and of the incorrect explanation treatment on the right.
4.3. AI Development
The AI used in the study was a large language model (LLM) that provided participants with a classification (AI advice) and an explanation for its reasoning, depending on the treatment they were assigned to. As LLM, we used Open AI’s GPT-4o model (model version 2024-05-13) through an Azure Open AI Studio instance. In the pre-phase of the study, we tested multiple prompt strategies in a workshop with three authors. The final prompts that were used are shown in Appendix A in Table 4. The LLM was prompted such that the correct explanation treatment provided an explanation that corresponded to its prediction, while the explanation did not match the prediction in the incorrect explanation treatment.
The definition of an incorrect explanation follows Morrison et al. (2024) in that the AI does not provide correct reasoning for its predicted class, independent of the ground truth of the image. For instance, if the AI made a correct prediction for an Art Nouveau building, but the explanation does not conform to this class, it is defined as incorrect. To avoid information overload (Arnold et al., 2023), we designed the explanations to be no longer than three sentences. As outlined in Section 3, we analyze the impact of incorrect explanations for cases where the AI classification is correct.
4.4. Recruitment
We recruited 186 participants from the United States through the platform Prolific.co and ran the study on August 12, 2024. Previous research indicates that this platform is a reliable source of research data (Peer et al., 2017; Palan and Schitter, 2018). Several screening mechanisms were implemented through the Prolific platform. With the filters, we targeted individuals who were fluent in the English language and had shown high quality in previous studies (100% completion rate). Our recruitment strategy was designed to not focus on participants with specific backgrounds, but admit participants without any further restrictions to be able to generalize our findings. Participants who met the stated criteria and completed the study’s requirements received a base payment of £. Additionally, we implemented an incentive structure: participants are incentivized to conduct the task correctly by providing a bonus for each correctly classified image. This should ensure that participants paid attention during the task and did not provide random answers. The bonus was 4 pennies for each correct answer and led to a potential maximum payment of £. As stated in our pre-registration, we excluded participants who did not finish the main task on time (within 30 minutes) or did not finish the entire study. We also excluded participants with obvious misbehavior (e.g., clicking through the cases and always providing the same answer). Additionally, we computed the overall mean and standard deviation across all treatments and winsorized at 2.5 SD above/below the mean. Applying this exclusion strategy, we ended up with 160 participants equally assigned to the four treatments (40 participants for each treatment).
4.5. Metrics
Similar to previous work (e.g., (Hemmer et al., 2023; Schoeffer et al., 2024)), we assessed participants’ task performance in classifying the architectural styles in pre-test, main task, and post-test. The task performance was used to approximate participants’ procedural knowledge in the pre-test and post-test and their human-AI team performance in the main task. As metrics for the task performance, we used accuracy and measured the ratio of correctly classified images over all images. Participants had to select one of the three different architectural styles for each image by selecting from a drop-down menu on each page of the task. This means that a random guess corresponded to 33.3% of performance. Aligning with Schoeffer et al. (2024), we also measured the correct adherence and detrimental overrides in the main task of the study.
We assessed participants’ reasoning ability for the three architectural styles through open-ended questions by asking them to describe and explain each style, thereby following the procedure of Chi et al. (1989). We rated each answer in terms of correctness by comparing it to the correct characteristics and features for each architectural style. By doing so, we were able to assess whether incorrect explanations impaired participants’ reasoning.
Finally, we established several control variables to investigate the potential underlying factors that might influence AI-assisted decision-making. In particular, we controlled for participants’ cognitive load as previous research suggests that the information in explanations displayed to humans can affect their decision-making behavior (Abdul et al., 2020; Hudon et al., 2021; Herm, 2023; Spitzer et al., 2024a). We measured participants’ cognitive load on a seven-point Likert scale by having them rate five validated items previous research has established (Senoner et al., 2024). In addition, we assessed participants’ AI trust, AI usefulness, and experience with AI by using items proven in previous research (Senoner et al., 2024), all of them also rated on a seven-point Likert scale. All items are shown in Table 5 in Appendix A.
5. Results
To address the research questions, we first conduct several statistical analyses to answer RQ: 1 in subsection 5.1. Subsequently, we address RQ: 2 and qualitatively assess the open-ended questionnaires in subsection 5.2. In the final subsection 5.3, we evaluate the impact of incorrect explanations on the human-AI team performance (RQ: 3).
5.1. RQ1: Impact of Incorrect Explanations on Procedural Knowledge
It took participants, on average, 14 minutes and 54 seconds to complete the study. Overall, 160 participants passed the attention check and finished the study according to the study protocol. Of these 160 participants, 79 were male, 77 were female, and four identified as diverse. To establish a baseline and ensure that all treatments began on an equal performance level in classifying the architectural styles, we conduct a one-way ANOVA on the pre-test performance scores across the four treatments (). These results fail to reject the null hypothesis, indicating no significant differences in pre-test performance among the four treatments. This finding suggests that participants across all treatments start with comparable levels of procedural knowledge to classify the architectural styles prior to the main task. This allows for a thorough interpretation of any differences observed in the post-test results, as they can be more readily attributed to the treatment effects rather than pre-existing differences.
The figure shows two subfigures. In subfigure A, the performance of participants across treatments is shown for the main task. In subfigure B, for the post-test.
Overall, the control group maintains the lowest performance at around , while the correct explanation treatment continues to lead with approximately accuracy. The AI classification and incorrect explanation treatments show similar post-test performances of about and , respectively.
To assess the significance of differences in post-test performance across treatments, we compare performance levels between treatments. A one-way ANOVA yields evidence of treatment effects (), indicating that the type of AI support in the main task influences participant performance in the post-test. To further examine these differences, we conduct pairwise comparisons using one-sided t-tests, assuming participants in the incorrect explanation treatment exhibit lower performance compared to participants in the other performance. We also report the corrected p-values according to the Holm-Bonferroni correction (Holm, 1979) (see Table 8 in Appendix A). There is no significant difference between the post-test performances in the incorrect explanation treatment and the control group (). Comparison with the AI treatment also shows no significant difference (). Notably, the comparison between incorrect and correct explanations yields a significant difference (), with incorrect explanations leading to inferior performance. With these findings, we find support for hypothesis 1. These results demonstrate that the accuracy of explanations plays a crucial role in AI-assisted decision-making, with incorrect explanations decreasing the task performance of humans in the post-test the most. Correct AI-generated explanations provide an additional performance advantage over incorrect explanations, underscoring the importance of explanation accuracy in AI-assisted tasks.
To account for potential confounding factors influencing the effect of the different AI supports on procedural knowledge, we run regression analyses with the post-test performance as dependent variables and model the type of AI support as the independent variable. We define cognitive load, AI trust, AI knowledge, and AI usefulness as confounding factors. The regression analysis examining the impact of different treatments on post-test performance reveals that the type of AI support significantly influences participants’ performance (see Table 9 in Appendix A). The model accounts for of the variance in post-test performance (). Specifically, participants in the correct explanation treatment exhibit the greatest improvement in post-test performance, with a significant positive effect (), followed by those in the AI treatment () and the incorrect explanation treatment (). The data supports the findings of the statistical tests. Interestingly, cognitive load is negatively associated with post-test performance (), suggesting that higher cognitive load impedes procedural knowledge.
Additionally, we analyze how the procedural knowledge develops from pre-test to post-test. We plot this difference in Figure 4. The figure illustrates whether procedural knowledge develops or degrades across different decision-making stages in the different treatments. Interestingly, the data reveals that the incorrect explanation treatment still leads to a procedural knowledge development. Conversely, the AI treatment and correct explanation treatment show the highest procedural knowledge development, exceeding the development in the incorrect explanation treatment.
The figure shows the difference in performance between the pre-test and post-test for the different treatments as bar plots. Additionally, the performance of the control group in pre-test and post-test is shown as line plots.
Furthermore, we run an ANOVA and Tukey HSD Post-Hoc Test which reveal significant differences in procedural knowledge gains across treatments (). Notably, the incorrect explanation treatment does not significantly differ from the AI treatment or the correct explanation treatment ( for both comparisons), but it shows a trend toward higher procedural knowledge gains compared to the control group (). The procedural knowledge development is significantly higher than in the control group (). This suggests that while incorrect explanations may increase the procedural knowledge in collaboration with an AI, the correctness of explanations reduces the effect compared to correct explanations.
5.2. RQ2: Impact of Incorrect Explanations on Reasoning
To analyze participants’ reasoning capabilities as they transition from knowing ”how” to understanding ”why”, we assess the answers they provided in the open-text questionnaire. The primary focus is on evaluating how the accuracy of AI-generated explanations influences the participants’ reasoning abilities to understand how incorrect explanations impact their knowledge.
We assess participants’ reasoning abilities following the procedure of Huang et al. (2024). By doing so, we use LLMs to evaluate the correctness of participants’ answers and compute a reasoning score: zero represents a completely incorrect or irrelevant answer, and five represents a completely correct and comprehensive answer. The scores are then analyzed using independent t-tests to compare the performance between treatments, and the Holm-Bonferroni method is used to correct them.
In the incorrect explanation treatment (), participants score lower on average than those in the control group (). However, the difference is not significant (). When comparing the incorrect explanation treatment to the AI treatment(), participants exposed to incorrect explanations perform significantly worse (). The comparison between the incorrect explanation treatment and the correct explanation treatment () shows similar results (). This suggests that incorrect explanations tend to decrease reasoning abilities compared to correct explanations or only AI advice. These results highlight the critical impact of explanation accuracy on reasoning, with incorrect explanations leading to notable decreases in reasoning performance, particularly when compared to AI-generated or correct explanations.
The figure shows bar plots for the different treatments regarding the reasoning scores calculated. Additionally, a dotted line is shown that represents the control group.
5.3. RQ3: Impact of Incorrect Explanations on Human-AI Team Performance
In this subsection, we analyze the data of the study to derive insights into how incorrect explanations affect the human-AI team performance.
The main task performance results reveal substantial differences across treatments (see Figure 3(b)). The control group demonstrates the lowest performance at . In contrast, all AI-assisted treatments show higher performance levels. The AI treatment achieves accuracy, while the correct explanation treatment performs best at roughly . Interestingly, the incorrect explanation treatment still outperforms the control group, reaching about accuracy. These patterns persist, albeit with reduced magnitudes, in the post-test performance (see Section 5.1. We can also see that with the different performance levels in each treatment, the correct adherence and the detrimental overrides change. Especially in the incorrect explanation treatment, the detrimental overrides are the highest, indicating that participants overrely on the AI. These results suggest that AI support, regardless of the explanation accuracy, enhances performance during the main task, with incorrect explanations leading to the lowest gain.
To assess the significance of differences in main task performance across treatments, we conduct a one-way ANOVA. The results () indicate significant differences among the treatments. To assess the impact of incorrect explanations with the other treatments on main task performance, we conduct pairwise comparisons between the incorrect explanation treatment and other treatments using one-sided t-tests, assuming that participants in the incorrect explanation treatment exhibit lower performance compared to participants in the other treatments (see hypothesis 3). We also correct the tests using the Holm-Bonferroni method and report these p-values (see Table 6 in Appendix A). The comparison between the incorrect explanation treatment and the control group shows no significant difference in performance (), indicating that the performance for incorrect explanations is not below the control group’s. Similarly, the comparison with the AI treatment yields no significant difference (). However, when comparing the incorrect explanation treatment with the correct explanation treatment, there is a significant difference in main task performance (), indicating that the incorrect explanations lead to a lower performance compared to the correct explanations. Thus, we find support for hypothesis 3.
Analog to the post-test performance, we run a regression analysis to reveal the effect of confounding factors on the main task performance (see Table 7 in Appendix A). The model investigating the main task performance explains of the variance (), showing that participants in AI-supported treatments outperform those in the control group. Participants in the correct explanation treatment demonstrate the most substantial improvement over the control group, with a significant positive effect (). This is closely followed by the AI treatment () and the incorrect explanation treatment (). Furthermore, AI trust shows a significant negative association with performance (), indicating a possible nuanced effect of trust on how participants interact with the AI.
These findings suggest that while incorrect explanations can aid immediate task performance, they may hinder the retention of procedural knowledge (as the post-test performance is below the main-task performance). We use a repeated measures approach and run a mixed-effects model to analyze how procedural knowledge is affected by first providing and then removing AI support. We define performance as the dependent variable, the type of AI support as the independent variable, the different stages—with AI support in the main task and without AI support in the post-test—as the mediating factor, cognitive load, AI trust, AI knowledge and AI usefulness as control factors and participant ID as a random factor.
The table shows the results of the mixed effects model with performance as the dependent variable. In the columns, the coefficient and the standard error are presented. As significance levels, .1, .05, and 0.01 are defined. Dependent variable Performance Coeff SE Intercept 0.695*** 0.084 AI support: - control group (baseline) - AI treatment 0.256*** 0.046 - correct explanation treatment 0.306*** 0.047 -incorrect explanation treatment 0.249*** 0.045 decision-making stage -0.096*** 0.035 AI treatment :decision-making stage -0.129*** 0.050 correct explanation treatment:decision-making stage -0.104** 0.050 incorrect explanation treatment:decision-making stage -0.131*** 0.050 cognitive_load -0.110 0.086 AI trust -0.187 0.132 AI knowledge -0.011 0.060 AI usefulness 0.063 0.095 participant_ID 0.016 0.029 Log-Likelihood 49.9483 Scale 0.0248
1
Note: * p ¡ .1; ** p ¡ .05; *** p ¡ .01
Participants in all AI treatments show significantly higher performance during the main task compared to the control group, with the correct explanation treatment leading to the highest performance (), followed by the AI treatment (), and the incorrect explanations (). However, when transitioning to the post-test, where AI support is removed, all AI-assisted treatments experience a significant decline in performance. The decrease is most pronounced for the incorrect explanation treatment (), indicating that while AI support with incorrect explanations initially boosts the performance during the main task, it negatively impacts the retention and development of procedural knowledge when the AI support is removed. The interaction effects in Figure 6(c) support this. The sub-figures illustrate that incorrect explanations can undermine procedural knowledge development, as evidenced by the greater performance drop in the post-test for participants who receive incorrect explanations (see Figure 6(c)) compared to those who receive correct explanations (see Figure 6(b)) or AI advice only (see Figure 6(a)) (lower slope in the lines for incorrect explanations).
The figure shows three sub-figures. In sub-figure A, the interaction effect of the transition from main task to post-test on the AI treatment, in figure B on the correct explanation treatment and in figure C on the incorrect explanation treatment.
6. Discussion
6.1. Summary of Findings
With the rise of AI in decision-making domains, it is crucial to understand how the interaction with AI affects decision-makers. Prior research so far has either focused on the implications of incorrect AI advice (e.g., (Schemmer et al., 2023c; Bansal et al., 2021)) or explored how the correctness of explanations affects trust and reliance on AI (e.g., (Morrison et al., 2024; Cabitza et al., 2024)). This work investigates how the correctness of AI explanations impacts humans through collaboration with an AI. In a classification task, participants were assigned to one of four treatments: no AI support, AI advice only, AI advice with correct explanations, and AI advice with incorrect explanations. We measured the task performance before, during, and post-collaboration to derive humans’ procedural knowledge, reasoning, and the human-AI team performance. The results show that all AI-assisted treatments led to significantly higher human-AI team performance during the main task compared to the control group, demonstrating the merits of AI support. However, the correctness of the explanations played a crucial role: Participants who had received correct explanations exhibited the highest main task performance, followed by those who received only AI advice without explanations, dominating those who had received incorrect explanations. Importantly, the findings reveal that the accuracy of the explanations has a lasting impact on procedural knowledge development. While incorrect explanations temporarily increase performance during the main task, the procedural knowledge was impaired post-collaboration, ultimately hindering knowledge retention compared to correct explanations or no explanations. Moreover, these results must be considered with caution. Even though there was an increase in procedural knowledge, participants’ reasoning capabilities in the incorrect explanation treatment were below the ones from the control group. This indicates that incorrect explanations can undermine the durable benefits of human-AI collaboration despite initial performance improvements. These results contribute to the HCI literature by underscoring the nuanced effects of explanation accuracy on human-AI collaboration. The findings emphasize the importance of designing AI-based support that provides accurate explanations to not only enhance human decision-making but also maintain and improve humans’ procedural knowledge and reasoning in the long run. Overall, the study’s key takeaway is that, while AI support can generally improve performance, the correctness of the explanations provided by the AI is a crucial determinant of human-AI collaboration, influencing humans’ ability to draw conclusions and perform the task autonomously post-collaboration with an AI. These insights can inform the design of more effective and human-centric AI-based decision support systems.
6.2. Implications
This work makes several contributions to the field of HCI by deepening the understanding of how the correctness of AI explanations influences humans’ procedural knowledge and reasoning ability, as well as the human-AI team performance. While previous research has often focused on the benefits of AI assistance in enhancing decision-making (Schemmer et al., 2023a), this study offers a more nuanced perspective, highlighting the critical role of explanation correctness on humans’ knowledge and understanding. It complements the large corpus of HCI literature on XAI (i.e.; (Morrison et al., 2024; Schoeffer et al., 2024; Schemmer et al., 2023c; Hemmer et al., 2023; Spitzer et al., 2024b)) by explicitly investigating the influence of incorrect explanations for scenarios when the AI advice is correct and identifies the negative repercussions for humans.
The misinformation effect of explanations. The study highlights a crucial risk associated with AI explanations: the potential misinformation effect to impair procedural knowledge and reasoning through incorrect explanations. The misinformation effect, extensively studied by Loftus et al. (1978); Loftus and Palmer (1974), describes how exposure to incorrect information can distort an individual’s memory of an event. This phenomenon extends to AI explanations, where incorrect explanations can similarly distort humans’ understanding. The integration of incorrect information through explanations into existing knowledge structures, as discussed by Ayers and Reder (1998), can lead to significant changes in both procedural knowledge and reasoning. Incorrect explanations, while potentially unharmful in the short term, can create an illusion of understanding, resulting in poorer performance in subsequent tasks without AI support. This phenomenon is particularly concerning in high-stakes environments such as healthcare or legal decision-making, where the quality of decision-making has far-reaching consequences (Topol, 2019). For organizations, this implies that the long-term efficacy of AI hinges not only on their immediate performance but also on their ability to foster accurate knowledge. Incorrect explanations can lead to a misalignment between the AI’s recommendations and humans’ understanding, which may diminish their overall effectiveness (Morrison et al., 2024). Yet, it also has the potential to impair humans post-collaboration with an AI. Therefore, organizations must prioritize the development of AI that provides correct and transparent explanations to ensure sustained, high-quality decision-making and prevent detrimental impacts on organizational knowledge and performance.
Taking a human-centric perspective. In our study, we could see that the human-AI team performance improved over the participants’ initial task performance. This pattern, also observed by prior research (Inkpen et al., 2023; Hemmer et al., 2024), showcases the potential of humans collaborating with AI. Even though the scenario in which the human-AI team performance exceeds the performance of human or AI alone—also referred to as complementary team performance (Bansal et al., 2019)—could not be reached (due to the study design choices this was not feasible as the AI advice was always correct), our findings showcase the two sides of explanations: while correct explanations improve the human-AI team performance compared to only receiving AI advice (bright side) and demonstrate the merits of XAI, incorrect explanations decrease the human-AI team performance compared to only receiving AI advice (dark side) and outline the pitfalls of XAI. Thus, it is important to implement mechanisms that allow humans to verify the correctness of explanations, for instance, through reflection mechanisms (Ehsan and Riedl, 2024). Furthermore, we could also identify AI trust as a confounding factor during the main task and cognitive load during the post-test. While these findings align with prior research (Ueno et al., 2022; Buçinca et al., 2021; Westphal et al., 2023), they also emphasize the important role of individuals’ characteristics and traits in human-AI collaboration. Therefore, research has to anticipate these factors in the design of robust and safe explanations.
The role of AI. While the focus of our research has mainly been on the humans in our AI-assisted decision-making study, the AI can take an important role in minimizing the risk of negative repercussions. In our study, these negative repercussions were expressed by humans’ decreasing procedural knowledge and reasoning capabilities post-collaboration. Although this decline was seen in every AI-supported treatment, incorrect explanations led to the greatest decline, with reasoning ability falling below that of the control group. Recent research in HCI should, therefore, focus on developing methods in which the AI itself can warn human collaborators of a potentially incorrect explanation, for instance, by using uncertainty scores (Huang et al., 2023; Prabhudesai et al., 2023) or cognitive forcing functions (Buçinca et al., 2021).
Anchoring on explanations. We also take a critical view on the phenomenon prior research has identified in collaboration scenarios between humans and AI in which the AI provides explanations: the anchoring effect on explanations (Wang et al., 2019; Bansal et al., 2021). The effect occurs when human decision-makers fixate on the explanation and form an incorrect understanding. In our study, participants in the incorrect explanation treatment demonstrated this behavior by achieving lower procedural knowledge on the post-test compared to the other AI-supported treatments and also by showing lower reasoning capabilities than participants in the control group. As a result, they made the right decisions for the wrong reasons, a condition often referred to as the Clever Hans Effect. (Schramowski et al., 2020; Kauffmann et al., 2020). Although research has adopted the term to describe AI behavior, in our study, humans demonstrate the effect. While this illustrates a major incision in the human knowledge structure, it is crucial for research and practice to develop mechanisms to counteract this effect.
Design guidelines for practice. Our findings confirm that AI support enhances immediate task performance, even when the explanations are incorrect. Even though this observation aligns with prior HCI studies demonstrating that explanations can increase human-AI team performance, such as those by Bansal et al. (2019) and Schemmer et al. (2023c), we extend prior research by revealing the negative impacts of explanations. Our work shows that the performance boost provided by AI during the collaboration can be short-term if the explanations are incorrect, as they negatively impact humans’ procedural knowledge and reasoning post-collaboration. This finding advances the existing literature on XAI, which has primarily emphasized factors like transparency, reliance, and user understanding during the collaboration(Lai and Tan, 2019; Ehsan et al., 2018; Morrison et al., 2024). Our research indicates that while correct explanations foster procedural knowledge development, incorrect explanations can lead to the erosion of knowledge over time, even when they temporarily enhance task performance in the collaboration with AI. The findings also underscore the importance of designing AI that prioritize the correctness of explanations, especially in contexts where AI support is not always provided and human decision-makers have to be able to decide autonomously, for instance, as required by the EU AI act (Commission, 2021). We suggest that AI should integrate mechanisms allowing humans to evaluate the correctness of explanations critically. For instance, incorporating counterfactual explanations, as explored by previous research (Wachter et al., 2017; Goyal et al., 2019; Schemmer et al., 2023b), could help users better understand the AI’s decision-making process and mitigate the adverse effects of incorrect explanations. Similarly, maintaining the human ability to perform tasks autonomously without AI support is favorable at the workplace. The role of XAI can impact the transfer of knowledge to counter risks of demographic change (Spitzer et al., 2024b). Incorrect explanations impairing knowledge development have the potential to endanger the successful business operations of organizations. Therefore, we advise designers to incorporate not only mechanisms to detect incorrect explanations post-hoc (Xu et al., 2023; Zhou et al., 2023; Hartwig et al., 2023) but also in the design of the AI.
The impact for organizations. Prior research has mainly explored effects that occur during the collaboration of humans and AI. For instance, Schemmer et al. (2023c) conceptualize the relationship between the appropriateness of reliance and how it relates to the human-AI team performance. Morrison et al. (2024) advance this view by the dimension of explanations and explore the effects of their correctness on humans’ appropriate AI reliance. By outlining potential downsides of XAI, this work addresses the impact of explanations’ correctness on humans post-collaboration. Our study demonstrates that their procedural knowledge and reasoning are impaired when they are provided with incorrect explanations and the AI support is removed. Taking in a human-centric perspective (Horvatić and Lipic, 2021), this repercussion presents harm to not only humans’ individual knowledge development (Bhatt, 2000) but also to their ability to provide meaningful assets to organizations (Nonaka and Takeuchi, 2007; Davenport, 1998). Maintaining individuals as valuable assets to organizations is crucial because it directly influences organizational innovation, efficiency, and adaptability in a competitive market. Sustaining humans’ knowledge development can foster a more resilient and informed workforce capable of driving sustained success (Grant, 1996).
By addressing these complex dynamics, this study contributes to the advancement of HCI as a field, offering practical insights for the design of AI that are both effective for human-AI collaboration and beneficial for humans’ procedural knowledge development.
6.3. Limitations and Future Work
Despite the valuable insights gained from this study, several limitations must be acknowledged, offering avenues for future research. First, the study explores how incorrect explanations affect humans in AI-assisted decision-making. To investigate the impact of such explanations, we designed a study with four treatments of different AI support: no support, only AI advice, AI advice with correct explanations, and AI advice with incorrect explanations. In real-world applications, the interaction with AI that provides only correct or incorrect explanations is rather unlikely. It presents valuable means to take the first steps to investigate the impact of incorrect explanations but does not reflect the real world. Future research could take on this aspect to extend our findings and evaluate how a mix of correct and incorrect explanations—a mix that is more realistic for deployed AI—affects humans’ procedural knowledge and reasoning. Especially, different ratios of the correctness of explanations could provide further insights and advance the field.
Second, the focus of the study was mainly on measuring the task performance to derive insights into humans’ procedural knowledge and inform about the impact on human-AI team performance. However, in real-world scenarios, performance might not be the only metric relevant. Other measures, like appropriate reliance (Schemmer et al., 2023c), trust or fairness (Schoeffer et al., 2024) in the AI, might also be of high relevance in AI-assisted decision-making as previous studies show (Hemmer et al., 2023; Bansal et al., 2021). It is of high importance to explore how these factors change over time and under the effect of incorrect explanations. Exploring the temporal implications can extend the views and offer new insights that support the robust and effective design of AI.
Lastly, the study primarily relied on short-term measures based on task performance, which may not fully capture the long-term impact of AI support on human knowledge and how the correctness of explanations impacts the human-AI team performance. Future research could employ longitudinal designs to assess how incorrect explanations influence procedural knowledge development and reasoning over extended periods and across multiple tasks. This approach would offer a deeper understanding of how different types of explanations contribute to sustained knowledge development, aligning with the principles of human-centered AI.
7. Conclusion
This work sets out the first steps towards investigating the effect of incorrect explanations on the human and the human-AI team. By doing so, we take a human-centric perspective and analyze the repercussions of incorrect explanations on task performance to derive insights into humans’ procedural knowledge and reasoning. In an online study, we assessed the impact of such explanations, specifically after the AI support is withdrawn, and humans must act autonomously.
With our work, we make several contributions to the HCI field: First, we identify a misinformation effect caused by incorrect explanations, which impairs humans’ procedural knowledge and reasoning. Second, we offer insights into how such incorrect explanations limit human-AI team capabilities. Finally, we provide guidelines for the effective and safe design of explanations that can foster AI-assisted decision-making. So we can eventually imagine: the AI provides a correct explanation for differentiating the architectural styles. You pass your exam.
Acknowledgements
Generative AI tools were utilized throughout this work. Specifically, ChatGPT, Claude and Github Copilot were employed to generate code for visualizations. Additionally, ChatGPT, DeepL Write, and Grammarly were used to enhance the writing quality of tutorials and explanations provided to participants during the experiments, as well as to improve the language across all sections of this paper.
Appendix A Appendix
The table shows the description of the three architecture styles used in the study. Architecture styles Descriptions Art Nouveau • Art Nouveau draws inspiration from the natural world, employing sinuous, sculptural, and organic shapes. The undulating asymmetry of these lines creates a sense of movement and vitality. • The style embraces arches and curving lines. Architects aimed to create something fresh and vital, resulting in buildings with unique and expressive forms. • Art Nouveau architects experimented with modern materials, including iron, glass, ceramics, and later concrete. The movement rejected rigid historical conventions in favor of innovation. Art Deco • Many Art Deco facades feature angular, upward-pointing vertical lines. These triangular shapes culminate in a series of steps, creating a sense of upward movement and dynamism. • Art Deco embraces bright, opulent colors. Stark contrasts, such as black and white or gold and silver, create a striking visual impact. • Art Deco buildings seamlessly blend stucco, terra-cotta, decorative glass, chrome, steel, and aluminum. These materials contribute to the sleek yet opulent appearance of structures. Georgian Architecture • Georgian buildings are meticulously symmetrical. Their facades exhibit a harmonious arrangement of windows, doors, and other architectural features. • Historic Georgian homes feature doors with fanlights above. Additionally, the doors are set back at least four inches from the brick face. • Georgian buildings favor time-tested materials. Bricks, stone, and wood form the backbone of these structures, emphasizing longevity and solidity.
The table shows the prompts used to generated the output of the LLM. Prompt for correct explanations Hello, you have to provide support to a human in classifying the architectural style of buildings in images. Overall, there are three different styles to choose from: Art Deco, Art Nouveau and Georgian Architecture. For an image, please also provide within two to three sentences an explanation for your decision. In the following is the image: Prompt for incorrect explanations Hello, you have to classify the architectural style of buildings in images. Overall, there are three different styles to choose from: Art Deco, Art Nouveau, and Georgian Architecture. Please provide the classification of an image. Additionally, we want to investigate how false explanations can mislead users. Thus, also explain your decision within 2-3 sentences. However, make sure that the explanation is incorrect, thus it does not correspond to your classification. Can you please provide a misleading explanation by using actual elements of the image? Here is the image:
The table shows the different variables and items used in the questionnaire of the study. Variable Items AI knowledge • How much experience do you have in working with an AI? AI usability • Using the AI system would enable others to classify architectural styles of buildings more quickly. • Using the AI system would improve the performance when classifying architectural styles of buildings. • Using the AI system would increase productivity for classifying the architectural styles of buildings. • Using the AI system would enhance my effectiveness of classifying the architectural styles of buildings. • Using the AI system would make it easier for classifying the architectural styles of buildings. • I would find the AI system useful for classifying the architectural styles of buildings. • Learning to operate the AI would be easy for me. • I would find it easy to get the AI to do what I want it to do. • My interaction with the AI would be clear and understandable. • I would find the AI to be flexible to interact with. • It would be easy for me to become skillful at using the AI. • I would find the AI easy to use. AI trust • The AI is deceptive. • I am suspicious of the AI’s intent, action, or outputs. • The AI’s actions will have a harmful outcome. • I am confident in the AI. • The AI is reliable. • I can trust the AI. Cognitive load • How mentally demanding was the task? • How physically demanding was the task? • How hurried or rushed was the pace of the task? • How hard did you have to work to accomplish your level of performance? • How insecure, discouraged, irritated, stressed, and annoyed were you?
The table shows the results of the statistical test for the main task to compare differences between the incorrect explanation treatment and the other treatments. Comparison t-statistic p-value p-value (corrected) incorrect explanations vs. control group 5.637 .999 1.000 incorrect explanations vs. AI -.546 .293 .293 incorrect explanations vs. correct explanations -2.359 .010 .021
The table shows the results of the regression model with the main task performance as the dependent variable. In the columns, the coefficient and the standard error are presented. As significance levels, .1, .05, and 0.01 are defined. Dependent variable Performance Coeff SE Intercept .676*** .083 condition_AI [True] .259*** .039 condition_correctexplanation [True] .311*** .040 condition_incorrectexplanation [True] .250*** .038 cognitive_load -.023 .087 AI trust -.237* .134 AI knowledge -.029 .061 AI attitude .092 .096 .367 Adj. .338 Log-Likelihood 59.892 F-statistic 12.580***
1
Note: * p ¡ .1; ** p ¡ .05; *** p ¡ .01
The table shows the results of the statistical test for the post-test to compare differences between the incorrect explanation treatment and the other treatments. Comparison t-statistic p-value p-value (corrected) incorrect explanations vs. control group 2.241 0.986 0.986 incorrect explanations vs. AI -0.371 0.356 0.356 incorrect explanations vs. correct explanations -1.742 0.043 0.085
The table shows the results of the regression model with the post-test performance as the dependent variable. In the columns, the coefficient and the standard error are presented. As significance levels, .1, .05, and 0.01 are defined. Dependent variable Performance Coeff SE Intercept .618*** .112 condition_AI [True] .125** .053 condition_correctexplanation [True] .198*** .054 condition_incorrectexplanation [True] .117** .052 cognitive_load -.197* .118 AI trust -.136 ,180 AI knowledge .006 .082 AI attitude .033 .130 .127 Adj. .087 Log-Likelihood 11.914 F-statistic 3.153***
1
Note: * p ¡ .1; ** p ¡ .05; *** p ¡ .01
The table shows the results of the ANOVA analysis to compare differences between the treatments for the difference in post-test performance and pre-test performance.
| Comparison | |||||
| Group 1 | Group 2 | Mean Difference | p-adj | Lower | Upper |
| AI | Control | -0.1667 | 0.0185 | -0.3129 | -0.0204 |
| AI | Correct Explanations | 0.0000 | 1.0000 | -0.1462 | 0.1462 |
| AI | Incorrect Explanations | -0.0250 | 0.9707 | -0.1712 | 0.1212 |
| Control | Correct Explanations | 0.1667 | 0.0185 | 0.0204 | 0.3129 |
| Control | Incorrect Explanations | 0.1417 | 0.0613 | -0.0046 | 0.2879 |
| Correct Explanations | Incorrect Explanations | -0.0250 | 0.9707 | -0.1712 | 0.1212 |